OpenStudy (anonymous):

Suppose that every person in the world (about 6 billion people in all) were employed as an atom counter. Each person would work a 40 hour week and be able to count one atom per second. How long would it take for 6 billion people to count the hydrogen atoms in 4.4 lb of hydrogen? Answer in units of years

OpenStudy (anonymous):

I know that 4.4 lbs of hydrogen atoms would weigh 1.201901904E27atoms.

OpenStudy (anonymous):

If it takes 1 second to weight an atom, you have to calculate first how many seconds would take to weight all the atoms. This is a simple rule of 3. 1 atom - 1 second 1.201901904E27atoms - 1.201901904E27 seconds Now we know the seconds it will take you have to calculate the weeks that will take to measure the atoms with all the people working 40 hours /week. Working seconds every week between all the people: \[Secs= 6\times10^{^{12}} people \times 40 hours \times \frac{ 60 \min }{ 1 hour } \times \frac{ 60 \sec }{ 1 \min } = 864000\times 10^{12} = 8.64\times 10^{17} seconds / week\] Now with another rule of 3: 8.64E17 sec -1 week 1.201901904E27 - ¿? \[\frac{ 1.201901904\times 10^{27} }{ 8.64 \times 10^{17} }=1.39109017\times 10^{9}\]

OpenStudy (anonymous):

1.3919017 E9 weeks of work.

OpenStudy (anonymous):

they want it in years not in weeks